### v1

# Optimal demonstration of generalized quantum contextuality

Abstract
Finding a set of empirical criteria fulfilled by any theory that satisfies the generalized notion of noncontextuality is a challenging task of both operational and foundational importance. The conventional approach of deriving facet inequalities from the relevant noncontextual polytope is computationally demanding. Specifically, the noncontextual polytope is a product of two polytopes, one for preparations and the other for measurements, and the dimension of the former typically increases polynomially with the number of measurements. This work presents an alternative methodology for constructing a polytope that encompasses the actual noncontextual polytope while ensuring that the dimension of the polytope associated with the preparations remains constant regardless of the number of measurements and their outcome size. In particular, the facet inequalities of this polytope serve as necessary conditions for noncontextuality. To demonstrate the efficacy of our methodology, we apply it to nine distinct contextuality scenarios involving four to nine preparations and two to three measurements to obtain the respective sets of facet inequalities. Additionally, we retrieve the maximum quantum violations of these inequalities. Our investigation uncovers many novel non-trivial noncontextuality inequalities and reveals intriguing aspects and applications of quantum contextual correlations.
I Introduction
The conflicting predictions of quantum theory with the notion of generalized noncontextuality, commonly known as generalized quantum contextuality, is a fundamental nonclassical feature of quantum theory Spekkens (2005); Liang et al. (2011); Mazurek et al. (2016); Schmid et al. (2018); Pusey (2018); Xu et al. (2016); Chaturvedi et al. (2021a); Budroni et al. (2022); Spekkens (2008); Chaturvedi and Saha (2020); Chaturvedi et al. (2021b); Pusey (2014); Lostaglio (2018); Schmid and Spekkens (2018); Schmid et al. (2020a, b); Catani et al. (2023a, 2022); Selby et al. (2022); Catani et al. (2023b); Lostaglio and Senno (2020); Tavakoli et al. (2021); Wright and Farkas (2023). Contextuality is equally significant from an operational perspective as it drives quantum-over-classical advantages in numerous information processing and communication tasks Spekkens et al. (2009); Chailloux et al. (2016); Saha and Chaturvedi (2019); Saha et al. (2019); Hameedi et al. (2017); Ghorai and Pan (2018); Ambainis et al. (2019); Schmid et al. (2022); Flatt et al. (2022); Roch i Carceller et al. (2022).
Noncontextuality adheres to the Leibnizian principle that assigns identical realist descriptions to operationally equivalent experimental procedures. For example, preparation noncontextuality assigns identical epistemic states to preparation procedures that are indistinguishable, meaning all measurements yield identical statistics. Similarly, measurement noncontextuality assigns identical response schemes to operationally indistinguishable measurement procedures. Generalized noncontextuality combines both preparation and measurement noncontextuality in scenarios involving prepare-and-measure experiments. Like Bell inequalities, generalized noncontextuality implies empirical inequalities known as noncontextuality inequalities (NCI). Quantum theory prescribes preparations and measurements that satisfy operational indistinguishability yet violate NCI.
Quantum theory can violate noncontextual inequalities involving nontrivial indistinguishability conditions among mixtures of preparation procedures, revealing preparation contextuality. Other notable nonclassical features, such as violations of Bell inequalities and the Kochen-Specker theorem, also implicate preparation contextuality within quantum theory. In addition to their foundational significance, remarkable applications of preparation contextuality have been proposed in arenas such as oblivious communication Spekkens et al. (2009); Chailloux et al. (2016); Saha et al. (2019), state discrimination Schmid and Spekkens (2018); Flatt et al. (2022), randomness certification Roch i Carceller et al. (2022).
A contextuality scenario is defined by the number of preparations, measurements, and measurement outcomes, along with the operational indistinguishability conditions for preparation and measurement procedures in their various convex mixtures. The work deals with a fundamental aspect of the study of contextuality. Given a contextuality scenario, finding a set of empirical criteria fulfilled by any noncontextual theory is a demanding task of both foundational and operational importance. As pointed out by Schmid et al. Schmid et al. (2018), in a contextuality scenario, the set of empirical statistics possessing noncontextual explanations forms a convex polytope. Consequently, the inequalities representing the facets of that polytope combine to provide the necessary and sufficient criteria for empirically witnessing noncontextuality. Schmid et al. Schmid et al. (2018) also formulate a computational technique to retrieve all the facet inequalities applicable to arbitrary contextuality scenarios. However, the method for determining the facets of the noncontextual polytope is computationally expensive. In particular, the noncontextual polytope is a product of two polytopes, one for preparations and the other for measurements. To obtain the facet inequality of the noncontextual polytope, one needs to compute the extremal points of a $D_{P}-$dimensional polytope associated with the preparations to find the extremal epistemic states, which are probability distributions over the ontic state space and the extremal points of a $D_{T}-$dimensional polytope associated with product polytope. It turns out, typically, $D_{P}$ increases polynomially with the number of measurements and $D_{T}$ increases polynomially with the number of measurements times the number of measurements, owing to the polynomial increase in the number of distinct ontic states one needs to consider. This feature of the noncontextual polytope differs strikingly from the respective polytopes of local correlations in Bell scenarios and noncontextual correlations in Kochen-Specker contextuality scenarios, wherein it suffices to consider a single ontic state. Therefore, it is highly desirable to seek efficient methods to find a set of empirical conditions, specifically statistical inequalities necessarily satisfied by noncontextual theories.
This work aims to address these fundamental aspects of contextuality. The computational challenge of retrieving all facet inequalities for arbitrary contextuality scenarios is significant. Unlike the polytopes of local correlations in Bell scenarios and noncontextual correlations in Kochen-Specker scenarios, which require only a single ontic state, the noncontextual polytope involves more complexity. Thus, it is desirable to find efficient methods to identify empirical conditions within the generalized noncontextuality framework. We propose a novel and efficient method for retrieving noncontextuality inequalities in any contextuality scenario, needing only a single ontic state to characterize the polytope for preparations. This approach, unlike conventional methods Schmid et al. (2018), maintains a constant polytope dimension regardless of the number of measurements and their outcomes. This allows us to obtain a polytope that includes the noncontextual polytope more quickly. The facet inequalities of this polytope are noncontextuality inequalities that any noncontextual theory must satisfy. Violations of these inequalities provide sufficient conditions for identifying generalized quantum contextual correlations.
Moreover, rigorous studies have so far been limited to only two contextuality scenarios: one with four preparations and two measurements with preparation indistinguishability conditions and another with six preparations and three measurements with indistinguishability conditions for both preparations and measurements Schmid et al. (2018); Pusey (2018). This work aims to provide a comprehensive analysis of other elementary contextuality scenarios and explore new applications of quantum contextuality in these scenarios.
The rest of the paper is organized as follows. We begin by providing an overview of the general scenario of contextuality. We then present our approach for obtaining the set of necessary conditions for noncontextuality in an arbitrary contextuality scenario. Next, we describe our method explicitly stating the algorithm for obtaining noncontextuality inequalities and discuss its merits in contrast to the standard approach. We investigate different contextuality scenarios using the proposed methodology by considering six scenarios with indistinguishability conditions only among preparations, as well as two additional scenarios with indistinguishability conditions among both preparations and measurements. These scenarios encompass a range of four to nine preparations and a maximum of three measurements. Consequently, we retrieve a large number of novel NCI as necessary conditions for noncontextuality in these scenarios. To obtain the maximum quantum violations of these NCI, we employ two semi-definite programming techniques introduced in Chaturvedi et al. (2021a). We also study the robustness with respect to the experimental noise of the quantum violations. Following that, we delve into various intriguing applications stemming from quantum violations of the discovered noncontextuality inequalities. We demonstrate that the quantum violation of some of these inequalities can serve to certify the dimensionality of quantum systems, non-projective measurements, and quantum randomness. Finally, we outline promising directions for future studies.
II Generalized notion of contextuality
Consider, a prepare and measure experiment entailing several distinct preparation and measurement procedures. A preparation procedure is labelled by $P_{x}$, where $x$ denotes the specific preparation, and a measurement procedure is denoted by $M_{z|y}$, where $z$ and $y$ represent the outcome and setting of the measurement, respectively. Using an operational theory, such as quantum theory, we can make predictions about the empirical statistics $\{p(z|x,y)\}$, where $p(z|x,y)$ indicates the probability of obtaining outcome $z$ when the measurement specified by $y$ is performed on the preparation specified by $x$. We say that two preparation procedures, $P_{x}$ and $P_{x^{\prime}}$, are operationally equivalent or indistinguishable (denoted as $P_{x}\sim P_{x^{\prime}}$) if they yield identical outcome statistics for all measurements,
$p(z|x,y)=p(z|x^{\prime},y),\ \forall M_{z|y}.$ (1)
Similarly, two measurement procedures $M_{z|y}$ and $M_{z^{\prime}|y^{\prime}}$ are operationally equivalent or indistinguishable (denoted as $M_{z|y}\sim M_{z^{\prime}|y^{\prime}}$), if they produce identical outcome statistics for all possible preparations
$p(z|x,y)=p(z^{\prime}|x,y^{\prime}),\ \forall P_{x}.$ (2)
Let us consider a prepare and measure experiment involving $n_{x}$ distinct preparations as $x\in\{0,\dots,n_{x}-1\}$ and $n_{y}$ different measurements as $y\in\{0,\dots,n_{y}-1\}$, with each measurement having $n_{z}$ possible outcomes as $z\in\{0,\dots,n_{z}-1\}$. In this experiment, we have a set of hypothetical preparations, each of which is realized by taking convex mixtures of these $n_{x}$ preparations such that the resultant mixed preparations are indistinguishable. These mixed preparations are labeled by the variable $s\in\{0,\dots,n_{s}\}$ and they are realized by the set of convex coefficients $\{\alpha_{x|s}\}$, satisfying
$\alpha_{x|s}\geqslant 0,\ \ \sum_{x}\alpha_{x|s}=1,\forall x,s.$ (3)
The indistinguishability conditions imply that for all $s,s^{\prime}\in\{0,\dots,n_{s}\}$,
$\sum_{x}\alpha_{x|s}P_{x}\sim\sum_{x}\alpha_{x|s^{\prime}}P_{x}.$ (4)
It is important to note that the above set of indistinguishability conditions are taken to be independent, meaning that no indistinguishability condition can be deduced from the other conditions. Mathematically, this requires the vectors $\{\vec{u}_{s}\}_{s}$ of these convex coefficients,
$\vec{u}_{s}=\left(\alpha_{0|s},\alpha_{1|s},\cdots,\alpha_{n_{x}-1|s}\right)$
to be an independent set of vectors.
Similarly, we have indistinguishable measurement procedures labelled by $t\in\{0,\dots,n_{t}\}$, each of them is realized by different taking convex mixtures with the coefficients $\{\beta_{z,y|t}\}$ satisfying
$\beta_{z,y|t}\geqslant 0,\ \sum_{z,y|t}\beta_{z,y|t}=1,\ \forall z,y,t.$ (5)
Thereupon, these indistinguishability conditions on measurements can be expressed as
$\sum_{z,y}\beta_{z,y|t}M_{z|y}\sim\sum_{z,y}\beta_{z,y|t^{\prime}}M_{z|y}\ ,$ (6)
for all $t,t^{\prime}\in\{0,\dots,n_{t}\}$. Here also, we consider these indistinguishability conditions to be independent, which requires the vectors
$\vec{v}_{t}=\left(\beta_{0|t},\beta_{1|t},\cdots,\beta_{n_{y}-1|t}\right)$
to form an independent set of vectors. The number of preparations, measurements, and measurement outcomes, together with the set of independent indistinguishability conditions, defines a contextuality scenario.
In quantum theory, preparations are described by density operators $\rho_{x}$, and measurements are described by positive semi-definite operators $\mathbbm{M}_{z|y}$, satisfying $\sum_{z}\mathbbm{M}_{z|y}=\mathbbm{1}$, where $\mathbbm{1}$ is the identity operator. The probability of obtaining outcome $z$ when performing measurement $M_{z|y}$ on preparation $P_{x}$ is given by $p(z|x,y)=\text{Tr}(\rho_{x}\mathbbm{M}_{z|y})$. Furthermore, quantum preparations and measurements satisfy the indistinguishability conditions given by (4) and (6) if and only if the following equalities hold:
$\sum_{x}\alpha_{x|s}\ \rho_{x}=\sum_{x}\alpha_{x|s^{\prime}}\ \rho_{x}\ ,\ \ \forall s,s^{\prime}\in\{0,\dots,n_{s}\}.$ (7)
and
$\sum_{z,y}\beta_{z,y|t}\mathbbm{M}_{z|y}=\sum_{z,y}\beta_{z,y|t^{\prime}}\mathbbm{M}_{z|y}\ ,\ \forall t,t^{\prime}\in\{0,\dots,n_{t}\}.$ (8)
An ontological model offers an explanation to the prediction of an operational theory by considering the state of the system to be an objective reality. This state is called the ontic state of the system, denoted by $\lambda\in\Lambda$, where $\Lambda$ is an arbitrary measurable space referred to as the ontic state space. A preparation procedure $P_{x}$ prepares the system in an ontic state $\lambda$ with probability $\mu(\lambda|x)$. The distribution $\mu(\lambda|x)$ is referred to as the epistemic state of the system. On the measurement side, the probability of obtaining the outcome $z$ when a measurement $M_{z|y}$ is performed on the ontic state $\lambda$ is given by the response function $\xi(z|\lambda,y)$. An ontological model satisfying the generalized notion of noncontextuality assigns identical epistemic states to indistinguishable mixed preparations and identical response functions to indistinguishable mixed measurement procedures Spekkens (2005). More precisely, the indistinguishability conditions (4) and (6) in any noncontextual ontological model imply
$\sum_{x}\alpha_{x|s}\mu(\lambda|x)=\sum_{x}\alpha_{x|s^{\prime}}\mu(\lambda|x),$ (9)
for all $s,s^{\prime}\in\{0,\dots,n_{s}\}$ and
$\sum_{z,y}\beta_{z,y|t}\xi(z|\lambda,y)=\sum_{z,y}\beta_{z,y|t^{\prime}}\xi(z|\lambda,y),$ (10)
for all $t,t^{\prime}\in\{0,\dots,n_{t}\}$, pertaining to every $\lambda$. An operational theory is called noncontextual if its predictions can be explained by a noncontextual ontological model. Given a contextuality scenario, the set of empirical probabilities $\{p(z|x,y)\}$ obtained from any noncontextual operational theory forms a polytope Schmid et al. (2018). A general form of a facet inequality of the noncontextual polytope is given by
$\sum_{x,y,z}c_{x,y,z}\ p(z|x,y)\leqslant\mathcal{C},$ (11)
where $c_{x,y,z}$ are real coefficients and $\mathcal{C}$ is the noncontextual bound. An operational theory whose predictions violate such inequalities is said to be contextual. There exists $\{p(z|x,y)\}$ predicted by quantum theory that violates such inequalities in general.
III Construction of the polytope and method to obtain necessary conditions for noncontextuality
This section presents our approach to derive NCI for any given contextuality scenario. The fundamental underpinning concept entails retrieving the extremal points of a polytope that characterizes the epistemic states corresponding to the preparations in an ontological model. Crucially, these extremal points are formulated to be independent of the number of ontic states while encompassing all feasible response functions describing the measurements.
Polytope characterizing preparations. Let us recall that the presence of indistinguishability conditions in a given contextuality scenario enforces the relationship (9) on the epistemic states ${\mu(\lambda|x)}$, where ontic state $\lambda$ can take arbitrary possible values. Let us now introduce the following additional variable,
$q(x,\lambda)=\frac{\mu(\lambda|x)}{\sum_{x}\alpha_{x|s}\mu(\lambda|x)}=:\frac{\mu(\lambda|x)}{\overline{\mu}(\lambda)}.$ (12)
Note that the denominator of the above expression is a constant for all $s$, and has the expression
$\overline{\mu}(\lambda)=\sum_{x}\alpha_{x|s}\mu(\lambda|x).$ (13)
By dividing both sides of Eq. (9) by $\overline{\mu}(\lambda)$, we find that our new variables $\{q(x,\lambda)\}$ satisfy the following conditions for all $\lambda$,
$\forall s,\ \sum_{x}\alpha_{x|s}q(x,\lambda)=1.$ (14)
Hence, for each $\lambda$, the collection of variables $\{q(x,\lambda)\}_{x}$ constitutes a convex polytope that adheres to the positivity constraint $q(x,\lambda)\geqslant 0$, as well as to the constraints implied by Eq. (14). Let’s denote the extremal points of this polytope as $e_{p}$, and the corresponding values at these extremal points as $q(x|e_{p})$. In simpler terms, any $q(x,\lambda)$ can be expressed in the following manner:
$q(x,\lambda)=\sum_{e_{p}}w(e_{p}|\lambda)q(x|e_{p}),$ (15)
where $w(e_{p}|\lambda)$ are convex weights relative to each specific $\lambda$.
Polytope characterizing measurements. On the other side, for each $\lambda$, the set of quantities $\{\xi(z|\lambda,y)\}_{z,y}$ forms a convex polytope that fulfills the criteria of positivity ($\xi(z|\lambda,y)\geqslant 0$), normalization ($\sum_{z}\xi(z|\lambda,y)=1$), and adheres to the constraints dictated by the indistinguishability condition in Eq. (10). We assign the label $e_{m}$ to the extremal points of this polytope, with the corresponding value at these points denoted by $\xi(z|y,e_{m})$. Consequently, any $\xi(z|\lambda,y)$ can be represented in the following way
$\xi(z|\lambda,y)=\sum_{e_{m}}w(e_{m}|\lambda)\xi(z|y,e_{m}).$ (16)
Here, $w(e_{m}|\lambda)$ constitutes a valid set of convex weights pertaining to each $\lambda$.
The combined polytope. We now articulate the empirical probability $p_{NC}(z|x,y)$ stemming from non-contextual operational theories in terms of the extremal distributions $q(x|e_{p})$ and $\xi(z|y,e_{m})$ using following sequence of mathematical relations:
[…]
$=$ $\int_{\lambda}\sum_{e_{m}}w(e_{m}|\lambda)\xi(z|y,e_{m})\mu(\lambda|x)\ d\lambda$
[…]
The second, fourth, and fifth lines can be deduced from their respective preceding lines through the application of Eqs. (16), (12), and (15), respectively. The third and sixth lines involve a rearrangement of the summation terms. At this juncture, we can introduce the variables
$w(e_{p},e_{m}|\lambda):=w(e_{p}|\lambda)w(e_{m}|\lambda),$ (18)
which form a set of convex weights since they satisfy $w(e_{p},e_{m}|\lambda)\geqslant 0$ and $\sum_{e_{p},e_{m}}w(e_{p},e_{m}|\lambda)=1$ for all $\lambda$. Additionally, by summing over $\lambda$ on both sides of Eq. (13), it is apparent that $\int_{\lambda}\overline{\mu}(\lambda)=1$. Hence, we have
$\int\overline{\mu}(\lambda)w(e_{p},e_{m}|\lambda)d\lambda\leqslant\max_{\lambda}w(e_{p},e_{m}|\lambda)=:w^{*}(e_{p},e_{m}),$ (19)
substituting this into (17), we arrive at the ensuing relation
$p_{NC}(z|x,y)\leqslant\sum_{e_{p},e_{m}}w^{*}(e_{p},e_{m})q(x|e_{p})\xi(z|y,e_{m}).$ (20)
Thus, the probabilities $p(z|x,y)$ in any non-contextual theory are necessarily bounded by and confined within the polytope whose extremal points emerge from the multiplication of $e_{p}$ and $e_{m}$. In other words, the probabilities $p(z|x,y)$ lie within the polytope determined by the extremal probability distributions $\{q(x|e_{p})\xi(z|y,e_{m})\}$. Let’s denote the probabilities associated with this bigger polytope as
$p(z|x,y)=\sum_{e_{p},e_{m}}w(e_{p},e_{m})q(x|e_{p})\xi(z|y,e_{m}).$ (21)
We can reconfirm, using the above relation, that $p(z|x,y)$ indeed satisfy the indistinguishability conditions (4) since
$\sum_{x}\alpha_{x|s}p(z|x,y)$ (22)
$=$ $\sum_{e_{p},e_{m}}w(e_{p},e_{m})\left(\sum_{x}\alpha_{x|s}q(x|e_{p})\right)\xi(z|y,e_{m})$
$=$ $\sum_{e_{m}}w(e_{m})\xi(z|y,e_{m})$
remains constant for all $s$, in accordance with (14). Similarly, $p(z|x,y)$ also adheres to the indistinguishability condition (6). However, the expression
$\sum_{z}p(z|x,y)=\sum_{e_{p}}w(e_{p})q(x|e_{p})$ (23)
does not necessarily equate to $1$, thus violating the normalization condition. Nevertheless, all observed probabilities should satisfy the normalization requirement:
$\sum_{z}p(z|x,y)=1.$ (24)
In fact, there exists a polytope wherein the probabilities conform solely to the normalization conditions (24). To rectify this issue with the extended polytope, our strategy is to identify the facet inequalities of the polytope defined by (21) while it is constrained by the polytope where probabilities adhere to the normalization conditions (24). These inequalities form NCI since the actual noncontextual polytope lies within both these polytopes. For a visual representation, refer to Fig. 1. Prior to presenting the succinct algorithm outlining our method to derive NCI, we first demonstrate the method by explicitly applying it to the simplest contextuality scenario.
Figure 1: Here, we denote P as the extended polytope encompassing the probabilities specified by (21); NP as the polytope that adheres solely to the normalization condition (24); and NCP as the precise non-contextual polytope, which exists within the confines of the other two polytopes. The noncontextuality inequalities (NCI) derived in this work are essentially the facet inequalities of the polytope formed by the intersection of P and NP.
An explicit example. The simplest contextuality scenario involves four preparations indexed by $x\in\{0,1,2,3\}$, such that obey the following indistinguishability condition
$\frac{1}{2}\left(P_{0}+P_{1}\right)\sim\frac{1}{2}\left(P_{2}+P_{3}\right).$ (25)
Furthermore, there are two measurements, indexed by $y\in{0,1}$, each yielding binary outcomes $z\in{0,1}$. No non-trivial indistinguishable conditions are applied to the measurements in this contextuality scenario. According to (14), the variables $\{q(x)\}$ describing the preparations adhere to the relations
$q(x)\geqslant 0,\ q(0)+q(1)=q(2)+q(3)=2,$ (26)
while the variables $\{\xi(z|y)\}$ characterizing the measurements satisfy
$\xi(z|y)\geqslant 0,\ \xi(0|0)+\xi(1|0)=\xi(0|1)+\xi(1|1)=1.$ (27)
Clearly, the set of allowed values of these variables form convex polytopes. It easy to see that there are four extremal points of both these polytopes which are labeled as $e_{p},e_{m}\in\{1,2,3,4\}$,
$e_{p}$ $q(0|e_{p})$ $q(1|e_{p})$ $q(2|e_{p})$ $q(3|e_{p})$
1 2 0 2 0
2 2 0 0 2
3 0 2 2 0
4 0 2 0 2
$e_{m}$ $\xi(0|0,e_{m})$ $\xi(1|0,e_{m})$ $\xi(0|1,e_{m})$ $\xi(1|1,e_{m})$
1 1 0 1 0
2 1 0 0 1
3 0 1 1 0
4 0 1 0 1
By combining these extremal points, we obtain $16$ extremal distributions $\{q(x|e_{p})\xi(z|y,e_{m})\}$, each corresponding to a specific pair $(e_{p},e_{m})$. Subsequently, we extract the facet inequalities of the polytope defined by these $16$ extremal distributions, resulting in a total of $24$ facet inequalities for this case. First, by utilizing the indistinguishable condition (25),
$\forall y,\ p(z|3,y)=p(z|0,y)+p(z|1,y)-p(z|2,y),$ (28)
we can substitute the probabilities $\{p(z|3,y)\}$ in the facet inequalities to make them simpler. Secondly, the polytopes of variables $\{q(x)\}$ and $\{\xi(z|y)\}$ satisfy symmetry conditions with respect to interchanging variables. For example, if variables $x=0$ and $x=1$ are interchanged, then the four extremal points $\{q(x|e_{p})\}$ remain unchanged, and consequently, the associated polytope retains the same form. Let us denote such symmetry by $P_{0}\longleftrightarrow P_{1}$. Similarly, we can identify all the symmetries in this scenario
$P_{0}\longleftrightarrow P_{1}$
$P_{2}\longleftrightarrow P_{3}$
$M_{0|y}\longleftrightarrow M_{1|y}\quad\forall y=0,1$
$M_{z|0}\longleftrightarrow M_{z|1}\quad\forall z=0,1.$ (29)
When an inequality becomes identical to another after applying any of the symmetries, it indicates that these two inequalities are equivalent. After systematically applying all of the symmetries (III), we can identify a set of inequalities that are equivalent to each other and form an equivalence class. It is sufficient to consider any one representative inequality from that equivalent set or equivalent class of inequalities. The ‘orbit size’ signifies the number of inequalities within a specific equivalence class. Among the $24$ inequalities initially obtained, the different in-equivalent classes of facet inequalities are enlisted below, where we have used a short-hand notation $p_{x,y}^{z}:=p(z|x,y)$.
orbit size Inequalities
9 $-p^{1}_{2,0}\leqslant 0$
3 $p^{0}_{0,0}+p^{0}_{1,0}+p^{1}_{2,0}\leqslant 2$
2 $-p^{0}_{0,0}-p^{1}_{0,0}+p^{1}_{0,1}\leqslant 0$
1 $p^{0}_{0,0}+p^{1}_{0,0}+p^{1}_{1,1}\leqslant 2$
3 $p^{0}_{1,0}+p^{1}_{2,0}+p^{1}_{0,1}-p^{1}_{2,1}\leqslant 2$
3 $-p^{0}_{2,1}-p^{1}_{0,0}+p^{1}_{2,0}-p^{1}_{1,1}\leqslant 0$
1 $-p^{0}_{0,0}-p^{0}_{1,0}+p^{0}_{2,1}-p^{1}_{2,0}+p^{1}_{2,1}\leqslant 0$
1 $p^{0}_{0,0}+p^{0}_{1,0}-p^{0}_{2,1}+p^{1}_{0,0}+p^{1}_{2,0}-p^{1}_{0,1}\leqslant 0$
1 $-p^{0}_{0,0}-p^{0}_{1,0}+p^{0}_{2,1}-p^{1}_{0,0}-p^{1}_{2,0}+p^{1}_{0,1}\leqslant 0$
Some of these inequalities are violated by quantum theory. It is apparent that the normalization condition does not hold since the quantity (23) is not $1$ for any of the extremal points of $e_{p}$ given above. Therefore, we will consider the facet inequalities of the polytope intersected with the polytope where the probabilities satisfy the normalization conditions (24). This can be readily done by substituting $p(1|x,y)$ in the facet inequalities using the normalization condition
$p(1|x,y)=1-p(0|x,y)\ \forall x,y.$ (30)
Further, imposing the symmetries (III), we find that the number of in-equivalent classes reduces to only two,
orbit size Inequalities
16 $p^{0}_{2,0}\leqslant 1$
8 $p^{0}_{1,0}-p^{0}_{2,0}-p^{0}_{0,1}+p^{0}_{2,1}\leqslant 1$
Table 1: We obtained $24$ inequalities for the simplest contextuality scenario described in (25). This set of inequalities reduces to two in-equivalent classes after applying the indistinguishable conditions, symmetries, and normalization conditions.
While the first inequality is trivial, the second inequality is violated in quantum theory, achieving the maximum value $\sqrt{2}\approx 1.414$. It is interesting to note that this second inequality is identical to the success metric of the parity oblivious random access codes Spekkens et al. (2009).
Algorithm to obtain the set of noncontextuality inequalities
Now, we present the algorithm, outlining the step-by-step process for deriving NCI for any prepare and measure contextuality scenario.
  • First obtain the extremal points $\{q(x|e_{p})\}$ of the polytope for the variables $\{q(x)\}$ satisfying the conditions:
    $(i)\ q(x)\geqslant 0;$ (31)
    $(ii)\ \forall s,\ \sum_{x}\alpha_{x|s}q(x)=1.$ (32)
  • Next, evaluate the extremal points $\{\xi(z|y,e_{m})\}$ of the polytope of the variables $\{\xi(z|y)\}$ satisfying the conditions:
    $(i)\ \xi(z|y)\geqslant 0;$ (33)
    $(ii)\ \forall y,\ \sum_{z}\xi(z|y)=1;$ (34)
    $(iii)\ \forall t,t^{\prime},\ \sum_{z,y}\beta_{z,y|t}\xi(z|y)=\sum_{z,y}\beta_{z,y|t^{\prime}}\xi(z|y).$ (35)
    In the case of only indistinguishability conditions on preparations, the polytope is obtained using conditions $(i)$ and $(ii)$ only.
  • Multiply the extremal points obtained in the previous steps to yield the extremal points of the extended polytope as ${q(x|e_{p})\cdot\xi(z|y,e_{m})}$. Subsequently, determine the facet inequalities of the extended polytope.
  • Impose the normalization condition (24) to the derived facet inequalities. For example, $p(z=k-1|x,y)$ can be replaced by $1-\sum_{z=0}^{k-2}p(z|x,y)$ for all $x,y$, so that the revised facet inequalities do not contain $p(k-1|x,y)$. Further reduce number of probabilities $\{p(z|x,y)\}$ in the facet inequalities using the indistinguishability conditions (4) and (6),
    $\forall s,s^{\prime},\forall z,y,\ \sum_{x}\alpha_{x|s}p(z|x,y)=\sum_{x}\alpha_{x|s^{\prime}}p(z|x,y),$ (36)
    $\forall t,t^{\prime},\forall x,\ \sum_{z,y}\beta_{z,y|t}p(z|x,y)=\sum_{z,y}\beta_{z,y|t^{\prime}}p(z|x,y).$ (37)
  • Finally, identify the symmetries on the variables $x,y,z,$ such that the extremal points $\{q(x|e_{p})\cdot\xi(z|y,e_{m})\}$ remain invariant, and apply these symmetries to the revised facet inequalities obtained in the previous step in order to obtain distinct in-equivalent classes of NCI.
Comparison with the method for finding the facets of exact noncontextual polyotpe
Method to find the exact noncontextual polytope was provided by Schmid et al. Schmid et al. (2018). The polytope presented in this work is notably larger than this exact noncontextual polytope. Consequently, the violation of the inequalities we derive here serves as a sufficient criterion (though not necessary) for operational certification of generalized contextuality. However, our approach presents a two-folded and substantial advantage over the method for identifying the exact noncontextual polytope in terms of efficiency.
We recall that $n_{x},n_{y},$ and $n_{z}$ refer to the number of preparations, measurements, and outcomes, respectively, in a contextuality scenario. Say, the number of extremal points obtained for the variables $\{\xi(z|y)\}$ satisfying (33)-(35) is $r$. According to the method in Schmid et al. (2018), the total number of ontic states $\lambda$ sufficient to characterize the preparations is $n_{x}\cdot r$. However, owing to the normalization conditions and the independent indistinguishability conditions (9), $n_{x}$ and $r\cdot n_{s}$ number of variables are eliminated, respectively. As a result, the dimension of the polytope characterizing the preparations becomes $(n_{x}-n_{s})r-n_{x}$ Henk et al. (2017).
In contrast, our method involves a fixed number of independent variables for characterizing the preparations, which is $n_{x}-n_{s}$ (see (31)-(32)) irrespective of the settings of the measurement side. Therefore, the difference between the dimensions of the two polytopes, whose extremal points are computed in the two different methods, is given by
$\Delta_{P}=(n_{x}-n_{s})r-2n_{x}+n_{s}.$ (38)
Furthermore, the method in Schmid et al. (2018) involves $r\cdot n_{y}\cdot n_{z}$ number of variables $\{\xi(z|y,\lambda)\}$ that describe the measurements. And, owing to the normalization conditions and the independent indistinguishability conditions (10), we can eliminate $r\cdot n_{y}$ and $r\cdot n_{t}$ number of variables, respectively. As a result, the dimension of the polytope characterizing the measurements becomes $r(n_{y}n_{z}-n_{y}-n_{t})$. One needs to compute the extremal points of the product of the two polytopes involving the variables $\{\mu(\lambda|x)\}$ and $\{\xi(z|y,\lambda)\}$ by multiplying the extremal points of these two polytopes. The product polytope has a dimension of $(n_{x}-n_{s})r-n_{x}+r(n_{y}n_{z}-n_{y}-n_{t})$, which follows from the fact that the dimension of a product polytope is the sum of the dimensions of the individual polytopes Henk et al. (2017). On the other hand, the product polytope, for which we compute the facet inequalities in our method, possesses a dimension of $n_{x}-n_{s}+n_{y}n_{z}-n_{y}-n_{t}$. Hence, the difference in dimensions between these two product polytopes, whose extremal points are computed through these two methods is given by
$\Delta_{T}=(r-1)(n_{x}+n_{y}n_{z}-n_{s}-n_{y}-n_{t})-n_{x}.$ (39)
As an example, when there are no indistinguishability conditions for measurements, we know that $r=n_{z}^{n_{y}}$ and $n_{t}=0$. Consequently, $\Delta_{P}$ scales in the order of $n_{z}^{n_{y}}$, and $\Delta_{T}$ scales in the order of $n_{y}n_{z}^{n_{y}}$, leading to a polynomial increase with the number of measurements.
IV Noncontextuality inequalities for various scenarios and their quantum violations
In this section, we explore various elementary contextuality scenarios in detail. First, we retrieve all the NCI using the aforementioned method. Subsequently, we conduct a comprehensive study of their quantum violations and establish upper bounds on the maximum quantum violations. Before presenting the results, we introduce a Fact that greatly aids in obtaining quantum violation for any preparation NCI with binary outcomes.
Fact 1.
Consider any linear figure of merit in any contextuality scenario where there are no nontrivial indistinguishability conditions on measurements, and the outcomes are binary, that is, $z\in\{0,1\}$. Since the outcomes are binary, we can replace $p(1|x,y)$ by $1-p(0|x,y)$ and express any linear expression as $\sum_{x,y}c_{x,y}p(0|x,y)$ where $c_{x,y}$ are some real numbers. Given any quantum preparation strategy $\{\rho_{x}\}$ satisfying the indistinguishability conditions, the best possible quantum measurement strategy $\{\mathbbm{M}_{z|y}\}$ is fully determined by the preparation strategy as
$\mathbbm{M}_{0|y}=\sum_{a>0}\mathbbm{P}_{a},$ (40)
where $\mathbbm{P}_{a}$ is the eigenprojector of the operator $\left(\sum_{x}c_{x,y}\rho_{x}\right)$ corresponds to positive eigenvalue $a$, that is, $\left(\sum_{x}c_{x,y}\rho_{x}\right)\mathbbm{P}_{a}=a\mathbbm{P}_{a}$.
Proof.
By splitting the sum and replacing the probabilities with quantum states and measurements, we find that
$\sum_{x,y}c_{x,y}p(0|x,y)$ $=$ $\sum_{y}\left(\sum_{x}c_{x,y}p(0|x,y)\right)$ (41)
$=$ $\sum_{y}\text{Tr}\left[\left(\sum_{x}c_{x,y}\rho_{x}\right)\mathbbm{M}_{0|y}\right].$
[…]
Two semi-definite programming-based methodsChaturvedi et al. (2021a) are used to identify quantum violations, with the see-saw method providing a lower bound on NCI violations. It generates random quantum states adhering to indistinguishability conditions on preparations.Quantum measurements are then optimized to maximize the relevant linear function of probabilities (the left-hand-side of the relevant NCI) while satisfying the indistinguishability conditions on measurements. In general, this set forms a semi-definite program, except in cases where there are no indistinguishability conditions on measurements and outcomes are binary, and Fact 1 is used to determine the optimal quantum measurements for the generated quantum states. The next step involves fixing optimized measurements and finding the best quantum states that optimize the expression while satisfying indistinguishability conditions on preparations. This semi-definite program is iterated until the expression’s value saturates. The optimization process is performed for different initial random quantum states, and the best value and strategy are retained. The value obtained through this see-saw method is denoted as $\mathcal{Q}^{d}_{s}$ for the chosen dimension ($d$) of the quantum states and measurements. While the value $\mathcal{Q}^{d}_{s}$ may not be the optimal quantum value, the method is useful as it delivers the corresponding quantum states and measurements that achieve this value.
Additionally, the robustness of the quantum violations is studied as a means of comparing violations of different NCI inequalities. Here, we consider the robustness with respect to the presence of white noise, which is the maximum amount of white noise that can be added while the quantum violation persists. Specifically, given the quantum states $\{\rho_{x}\}_{x}$ and quantum measurements $\{\mathbbm{M}_{z|y}\}_{z,y}$ acting on $\mathbbm{C}^{d}$ achieving $\mathcal{Q}^{d}_{s}$ found from the see-saw method, we take the noisy states $\omega(\mathbbm{1}/d)+(1-\omega)\rho_{x}$, where $\omega\in[0,1]$ being the noise parameter. As the measure of robustness, the minimum value of $\omega$, denoted by $\omega^{d}_{c}$, is then determined so that the left-hand side of NCI coincides with the noncontextual bound $\mathcal{C}$. It can be readily verified that, for a general inequality (11),
$\omega^{d}_{c}=\frac{\mathcal{Q}^{d}_{s}-\mathcal{C}}{\mathcal{Q}^{d}_{s}-\gamma},$ (42)
where $\gamma=(1/d)\sum_{x,y,z}c_{x,y,z}\text{Tr}(\mathbbm{M}_{z|y})$ is the value of left-hand-side of (11) for the maximally mixed state. The larger value of $\omega^{d}_{c}$ the more robust is the quantum violation arising from $\{\rho_{x}\}_{x}$ and $\{\mathbbm{M}_{z|y}\}_{z,y}$.
The second method involves implementing the semi-definite hierarchy introduced in Chaturvedi et al. (2021b) up to the first level, which yields a dimension independent upper bound on the maximum quantum violation of NCI. We denote this upper bound by $\mathcal{Q}_{1}$. Therefore, if $\mathcal{Q}_{1}$ matches with $\mathcal{Q}^{d}_{s}$, it indicates the exact maximum violation (up to machine precision).
The simplest scenario of four preparations satisfying indistinguishability conditions (25) has been discussed rigorously in Sec. III. Eight other scenarios with their quantum violations are discussed in the following subsections. The respective polytopes and symmetries have been found using ‘polymake’, and the bounds on the quantum violations based on the aforementioned semi-definite programming methods are found using ‘Matlab’ and ‘sdpt3’. The codes are available in cod (2024). We enlist the NCI except the trivial NCI that are of the form $p(z|x,y)\leqslant 1$ or $p(z|x,y)\geqslant 0$. Apart from the last scenario, all other scenarios involve only binary outcomes. For the sake of convenience thereof, we express the NCI only in terms of outcome $z=0$ and further use the following short-hand notation to denote the probabilities
$p_{x,y}:=p(0|x,y).$ (43)
Scenario 2
The contextuality scenario is defined as follows:
$x\in\{0,1,2,3\},\ y\in\{0,1,2\},\ z\in\{0,1\},\quad\frac{1}{3}\left(P_{0}+P_{1}+P_{2}\right)\sim\frac{1}{2}\left(P_{0}+P_{3}\right).$ (44)
orbit size Inequalities $\mathcal{Q}_{s}^{2}$ $\omega^{2}_{c}$ $\mathcal{Q}_{1}$
24 $\mathcal{I}_{2}=-p_{0,0}+2p_{1,0}+p_{0,1}-2p_{2,1}\leqslant 2$ 2.6458 0.244 2.7321
5 $p_{0,2}-2p_{1,2}-2p_{2,2}\leqslant 0$ 0 0 0
Table 2: We obtained $48$ inequalities in the contextuality scenario (44). Out of these inequalities, $19$ are trivial. The rest of them are reduced to $2$ inequivalent classes after applying the indistinguishable conditions and symmetries mentioned below.
$\forall y,z,\ p(z|x=3,y)=\frac{2}{3}p(z|x=1,y)+\frac{2}{3}p(z|x=2,y)-\frac{1}{3}p(z|x=0,y);$ (45)
$P_{1}\longleftrightarrow P_{2};\ M_{0|y}\longleftrightarrow M_{1|y}\ \forall y;\ \ M_{z|0}\longleftrightarrow M_{z|1};\ \ M_{z|1}\longleftrightarrow M_{z|2};\ M_{z|0}\longleftrightarrow M_{z|2}\forall z.$ (46)
The set of inequivalent NCI is enlisted in Table 46, and the quantum strategy that achieves the highest quantum violation of $\mathcal{I}_{2}$ is illustrated in Fig. 2. The set of inequivalent nontrivial NCI is listed in Tab. 46. An interesting aspect of this contextuality scenario deserves attention. We define an ontic distribution as deterministic when $\mu(\lambda|x)\in{0,1}$; otherwise, it’s probabilistic. Notably, the noncontextual bound 2 of the NCI $\mathcal{I}_{2}$ (in Tab.46) can only be attained through a probabilistic ontic distribution in any ontological model. For instance, consider the following model where $\Lambda\in{\lambda_{1},\lambda_{2}}$:
$\frac{1}{3}(\mu(\lambda|0)+\mu(\lambda|1)+\mu(\lambda|2))=\frac{1}{2}(\mu(\lambda|0)+\mu(\lambda|3)).$ (47)
We say that an epistemic state is deterministic if $\mu(\lambda|x)\in\{0,1\}$; otherwise, it is probabilistic. The noncontextual bound $2$ can only be achieved by probabilistic epistemic state in any ontological model. For example, take the following model, where $\Lambda\in\{\lambda_{1},\lambda_{2}\}$
$x$ 0 1 2 3
$\mu(\lambda_{1}|x)$ 0 1 0 $\frac{2}{3}$
$\mu(\lambda_{2}|x)$ 1 0 1 $\frac{1}{3}$
$p_{d}(0|0,\lambda_{1})$ 1
$p_{d}(0|0,\lambda_{2})$ 0
$p_{d}(0|1,\lambda_{2})$ 0
This model satisfies (47) while simultaneously yielding $\mathcal{I}_{2}=2$. On the other hand, to determine the maximum value of $\mathcal{I}_{2}$ for deterministic epistemic states, it is sufficient to examine all feasible deterministic epistemic states considering at most four distinct ontic states. The optimization reveals that the value is $1$.
Scenario 3
The contextuality scenario consists of five preparations and two binary outcome measurements and is defined as follows:
$x\in\{0,1,2,3,4\},\ y\in\{0,1\},\ z\in\{0,1\},\quad\frac{1}{2}\left(P_{0}+P_{1}\right)\sim\frac{1}{3}\left(P_{2}+P_{3}+P_{4}\right).$ (48)
The set of inequivalent NCI for the contextuality scenario is given in Table 3, and the strategy that attains maximum quantum violation of $\mathcal{I}_{3}$ is illustrated in Fig. 2.
orbit size Inequalities $Q_{s}^{2}$ $\omega^{2}_{c}$ $Q_{1}$
2 $-3p_{0,0}-3p_{1,0}+2p_{2,0}+2p_{3,0}\leqslant 0$ 0 0 00
24 $\mathcal{I}_{3}=-3p_{0,0}+2p_{2,0}-3p_{1,1}+2p_{2,1}\leqslant 2$ 3.1231 0.272 $3.1815$
Table 3: We obtained $44$ inequalities in this scenario, among which $18$ are trivial. The rest of the inequalities are reduced to $2$ inequivalent classes after employing the following indistinguishability conditions and symmetries.
$\ p(z|x=4,y)$ $=$ $\frac{3}{2}\left(p(z|x=0,y)+p(z|x=1,y)\right)-p(z|x=2,y)-p(z|x=3,y)\ \forall y,z$ (49)
$P_{0}\longleftrightarrow P_{1};P_{2}\longleftrightarrow P_{3};P_{2}\longleftrightarrow P_{4};P_{3}\longleftrightarrow P_{4}$ (50)
$\ M_{0|y}\longleftrightarrow M_{1|y}\ \forall y;M_{z|0}\longleftrightarrow M_{z|1}\ \forall z$ (51)
$\ M_{0|0}\longrightarrow M_{0|1}\ ,\ M_{0|1}\longrightarrow M_{1|0}\ ,\ M_{1|0}\longrightarrow M_{1|1}\ ,\ M_{1|1}\longrightarrow M_{0|0}.$ (52)
[…]
Scenario 4
Table 57 contains the inequivalent NCI for the contextuality scenario defined as follows:
$x\in\{0,1,2,3,4\},y\in\{0,1\},z\in\{0,1\}\quad\frac{1}{4}\left(P_{0}+P_{1}+P_{2}+P_{3}\right)\sim\frac{1}{3}\left(P_{0}+P_{1}+P_{4}\right).$ (53)
orbit size Inequalities $Q_{s}^{2}$ $\omega^{2}_{c}$ $Q_{1}$
4 $p_{0,0}+p_{1,0}-3p_{2,0}-3p_{3,0}\leqslant 0$ 0 0 0
8 $p_{0,0}+p_{1,0}-3p_{2,0}+p_{0,1}+p_{1,1}-3p_{3,1}\leqslant 2$ 3.1231 0.272 $3.1815$
16 $p_{1,0}-3p_{2,0}+p_{1,1}-3p_{3,1}\leqslant 1$ 1.7417 0.198 $1.9168$
Table 4: We obtained $44$ inequalities including 16 trivial ones. The 28 nontrivial NCI are reduced to 3 inequivalent classes after applying the following indistinguishability conditions and symmetries.
$\ p(z|x=4,y)=\frac{3}{4}(p(z|x=2,y)+p(z|x=3,y))-\frac{1}{4}(p(z|x=0,y)+p(z|x=1,y))\ \forall y,z$ (54)
$P_{0}\longleftrightarrow P_{1};P_{2}\longleftrightarrow P_{3};$ (55)
$\ M_{0|y}\longleftrightarrow M_{1|y}\ \forall y;\ \ M_{z|0}\longleftrightarrow M_{z|1}\ \forall z$ (56)
$\ M_{0|0}\longrightarrow M_{0|1}\ ,\ M_{0|1}\longrightarrow M_{1|0}\ ,\ M_{1|0}\longrightarrow M_{1|1}\ ,\ M_{1|1}\longrightarrow M_{0|0}.$ (57)
Scenario 5
The contextuality scenario consisting of six preparations and three binary outcome measurements is defined as follows:
$x\in\{0,1,2,3,4,5\},y\in\{0,1,2\},z\in\{0,1\},\quad\frac{1}{2}\left(P_{0}+P_{1}\right)\sim\frac{1}{2}\left(P_{2}+P_{3}\right)\sim\frac{1}{2}\left(P_{4}+P_{5}\right).$ (58)
The set of inequivalent NCI is given in Table 62, and the qubit strategy that achieves the maximum quantum violation of $\mathcal{I}_{5}$ is provided in Fig. 2.
orbit size Inequalities $Q_{s}^{2}$ $\omega^{2}_{c}$ $Q_{1}$
24 $p_{0,0}+p_{1,0}-p_{4,0}\leqslant 1$ 1 0 1
24 $p_{0,0}-p_{2,0}-p_{1,2}+p_{2,2}\leqslant 1$ 1.4142 0.293 1.4142
24 $p_{1,0}-p_{4,0}+p_{0,2}-p_{4,2}\leqslant 1$ 1.4142 0.293 1.4142
24 $p_{2,1}-p_{4,1}-p_{0,2}-p_{1,2}+p_{2,2}+p_{4,2}\leqslant 1$ 1.4142 0.293 1.4142
192 $\mathcal{I}_{5}=p_{2,0}-p_{4,0}+p_{1,1}-p_{2,1}-p_{4,1}-p_{0,2}+p_{2,2}+p_{4,2}\leqslant 2$ 2.5 0.2 2.6571
192 $-p_{0,0}+p_{2,0}+p_{4,0}+p_{1,1}-p_{2,1}-p_{1,2}-p_{2,2}+p_{4,2}\leqslant 2$ 2.5 0.2 2.6571
192 $p_{1,0}-p_{4,0}+p_{0,1}+p_{2,1}-p_{4,1}-2p_{0,2}-p_{1,2}+p_{2,2}+p_{4,2}\leqslant 2$ 2.5 0.2 2.6571
Table 5: Here we obtained 684 inequalities. Among these, 12 are trivial. The remaining inequalities are reduced to $8$ inequivalent classes after applying the following indistinguishability conditions and symmetry transformations.
$\ p(z|x=3,y)=p(z|x=0,y)+p(z|x=1,y)-p(z|x=2,y)\ \forall y,z$ (59)
$\ p(z|x=5,y)=p(z|x=0,y)+p(z|x=1,y)-p(z|x=4,y)\ \forall y,z$ (60)
$P_{0}\longleftrightarrow P_{1};P_{2}\longleftrightarrow P_{3};P_{4}\longleftrightarrow P_{5}$ (61)
$\ M_{0|y}\longleftrightarrow M_{1|y}\ \forall y;\ M_{z|0}\longleftrightarrow M_{z|1};\ M_{z|1}\longleftrightarrow M_{z|2};\ M_{z|0}\longleftrightarrow M_{z|2}\ \forall z.$ (62)
Scenario 6
The contextuality scenario consisting of seven preparations and three measurements is defined as follows:
$x\in\{0,1,2,3,4,5,6\},y\in\{0,1,2\},z\in\{0,1\},\quad\frac{1}{4}\left(P_{0}+P_{1}+P_{2}+P_{3}\right)\sim\frac{1}{3}\left(P_{4}+P_{5}+P_{6}\right).$ (63)
Table 68 presents the collection of inequivalent NCI obtained in this scenario. In contrast to previous scenarios, qubit strategy in this context did not yield the optimal results (using the see-saw method). In particular, there is a substantial enhancement in $Q_{s}$ when considering states and measurements of higher dimensions. We have documented these values up to $d=4$. Notably, for $\mathcal{I}_{6}^{2}$ and $\mathcal{I}_{6}^{4}$, we obtain $Q^{7}_{s}=14.9711$ and $Q_{s}^{7}=15.6163$, respectively, which are more than $Q_{s}^{4}$ given in table 68. In Fig. 2, we provide the details of the qubit strategy which violates $\mathcal{I}_{6}^{3}$. Here, we present the $3$-dimensional quantum states and measurements that yield a quantum violation of $\mathcal{I}_{6}^{1}=8.7764>6$:
$\rho_{0}=\Bigg{(}\begin{smallmatrix}0.7865&-0.1091&0.3950\\ -0.1091&0.0151&-0.0548\\ 0.3950&-0.0548&0.1983\end{smallmatrix}\Bigg{)},\rho_{1}=\Bigg{(}\begin{smallmatrix}0.0004&0.0186&0.0077\\ 0.0186&0.8526&0.3540\\ 0.0077&0.3540&0.1470\end{smallmatrix}\Bigg{)},\rho_{2}=\Bigg{(}\begin{smallmatrix}0.0004&0.0186&0.0077\\ 0.0186&0.8526&0.3540\\ 0.0077&0.3540&0.1470\end{smallmatrix}\Bigg{)},\rho_{3}=\Bigg{(}\begin{smallmatrix}0.3904&0.2225&-0.4341\\ 0.2225&0.1268&-0.2474\\ -0.4341&-0.2474&0.4828\end{smallmatrix}\Bigg{)}$
$\rho_{4}=\Bigg{(}\begin{smallmatrix}0.1713&0.2578&0.1378\\ 0.2578&0.5426&0.0428\\ 0.1378&0.0428&0.2860\end{smallmatrix}\Bigg{)},\rho_{5}=\Bigg{(}\begin{smallmatrix}0.5406&-0.4028&-0.2935\\ -0.4028&0.3000&0.2187\\ -0.2935&0.2187&0.1593\end{smallmatrix}\bigg{)},\rho_{6}=\Bigg{(}\begin{smallmatrix}0.1713&0.2579&0.1378\\ 0.2579&0.5427&0.0429\\ 0.1378&0.0429&0.2860\end{smallmatrix}\Bigg{)},$
$\mathbbm{M}_{0|0}=\Bigg{(}\begin{smallmatrix}0.8633&-0.3370&0.0668\\ -0.3370&0.1315&-0.0261\\ 0.0668&-0.0261&0.0052\end{smallmatrix}\Bigg{)}\ \mathbbm{M}_{0|1}=\Bigg{(}\begin{smallmatrix}0.5859&-0.0485&-0.4902\\ -0.0485&0.0040&0.0406\\ -0.4902&0.0406&0.4101\end{smallmatrix}\Bigg{)},\ \mathbbm{M}_{0|2}=\Bigg{(}\begin{smallmatrix}0.7889&0.3566&0.1984\\ 0.3566&0.3976&-0.3351\\ 0.1984&-0.3351&0.8136\end{smallmatrix}\Bigg{)}.$ (64)
Moreover, the following set of $4-$dimensional quantum states and measurements achieve the quantum violation $\mathcal{Q}_{s}=15.5037>12$ of $\mathcal{I}_{6}^{4}$ in Tab. 68,
$\rho_{0}=\Bigg{(}\begin{smallmatrix}0.1537&-0.1618&-0.1204&-0.2990\\ -0.1618&0.1704&0.1268&0.3148\\ -0.1204&0.1268&0.0943&0.2342\\ -0.2990&0.3148&0.2342&0.5816\end{smallmatrix}\Bigg{)},\rho_{1}=\Bigg{(}\begin{smallmatrix}0.3830&-0.3343&0.0919&0.3408\\ -0.3343&0.2917&-0.0802&-0.2974\\ 0.0919&-0.0802&0.0220&0.0818\\ 0.3408&-0.2974&0.0818&0.3032\end{smallmatrix}\Bigg{)},\rho_{2}=\Bigg{(}\begin{smallmatrix}0.1411&0.2127&-0.1319&0.0790\\ 0.2127&0.4379&-0.0719&0.0026\\ -0.1319&-0.0719&0.2609&-0.2001\\ 0.0790&0.0026&-0.2001&0.1601\\ \end{smallmatrix}\Bigg{)}$
$\rho_{3}=\Bigg{(}\begin{smallmatrix}0.1411&0.2127&-0.1319&0.0790\\ 0.2127&0.4379&-0.0718&0.0026\\ -0.1319&-0.0718&0.2609&-0.2001\\ 0.0790&0.0026&-0.2001&0.1601\end{smallmatrix}\Bigg{)},\rho_{4}=\Bigg{(}\begin{smallmatrix}0.1570&-0.2543&-0.0758&0.2489\\ -0.2543&0.4119&0.1227&-0.4031\\ -0.0758&0.1227&0.0366&-0.1201\\ 0.2489&-0.4031&-0.1201&0.3945\end{smallmatrix}\Bigg{)},\rho_{5}=\Bigg{(}\begin{smallmatrix}0.0102&0.0748&-0.0141&0.0657\\ 0.0748&0.5477&-0.1035&0.4811\\ -0.0141&-0.1035&0.0196&-0.0909\\ 0.0657&0.4811&-0.0909&0.4225\end{smallmatrix}\bigg{)}$
$\rho_{6}=\Bigg{(}\begin{smallmatrix}0.4470&0.1266&-0.1294&-0.1646\\ 0.1266&0.0438&-0.0921&-0.0610\\ -0.1294&-0.0921&0.4225&0.1478\\ -0.1646&-0.0610&0.1478&0.0867\end{smallmatrix}\Bigg{)},\ \mathbbm{M}_{0|0}=\Bigg{(}\begin{smallmatrix}0.0406&-0.1190&-0.0263&-0.1552\\ -0.1190&0.3488&0.0772&0.4550\\ -0.0263&0.0772&0.0171&0.1007\\ -0.1552&0.4550&0.1007&0.5935\end{smallmatrix}\Bigg{)}$
$\mathbbm{M}_{0|1}=\Bigg{(}\begin{smallmatrix}0.1083&0.1665&-0.1530&0.2133\\ 0.1665&0.9667&0.0656&-0.0115\\ -0.1530&0.0656&0.3433&-0.4447\\ 0.2133&-0.0115&-0.4447&0.5816\end{smallmatrix}\Bigg{)},\ \mathbbm{M}_{0|2}=\Bigg{(}\begin{smallmatrix}0.3099&-0.3260&0.0207&0.3274\\ -0.3260&0.3429&-0.0217&-0.3444\\ 0.0207&-0.0217&0.0014&0.0218\\ 0.3274&-0.3444&0.0218&0.3459\end{smallmatrix}\Bigg{)}.$ (65)
orbit Inequalities $Q_{s}^{2}$ $\omega_{c}^{2}$ $Q_{s}^{3}$ $\omega_{c}^{3}$ $Q_{s}^{4}$ $\omega_{c}^{4}$ $Q_{1}$ $Q^{2}_{UB}$
size
4 $3p_{0,0}+3p_{1,0}+3p_{2,0}+3p_{3,0}-4p_{4,0}-4p_{5,0}\leqslant 4$ 4 0 4 0 4 0 4 4
864 $\mathcal{I}_{6}^{1}=-3p_{1,0}-3p_{2,0}-3p_{3,0}+4p_{5,0}-3p_{0,1}-3p_{1,1}$
$-3p_{2,1}+4p_{5,1}+3p_{0,2}+3p_{3,2}-4p_{5,2}\leqslant 6$ 7.6833 0.144 8.7764 0.257 8.7764 0.270 9.2621 7.6833
864 $\mathcal{I}_{6}^{2}=-6p_{0,0}-3p_{2,0}-3p_{3,0}+4p_{4,0}+8p_{5,0}$
$-3p_{0,1}-3p_{1,1}+4p_{4,1}-6p_{1,2}-3p_{2,2}$ 12 0 14.4414 0.161 14.6929 0.171 16.9615 12
$-3p_{3,2}+4p_{4,2}+8p_{5,2}\leqslant 12$
864 $-3p_{3,0}+4p_{4,0}-3p_{0,1}-3p_{2,1}$
$+4p_{4,1}+3p_{1,2}-4p_{4,2}\leqslant 7$ 8.9692 0.197 8.9692 0.197 9.472 0.236 10.5888 8.9692
288 $-3p_{1,0}+4p_{5,0}+3p_{0,2}+3p_{2,2}+3p_{3,2}-4p_{5,2}\leqslant 9$ 10.8281 0.233 10.8281 0.267 10.8281 0.277 11.1229 10.8281
216 $\mathcal{I}_{6}^{3}=-3p_{1,0}-3p_{3,0}+4p_{5,0}+3p_{0,2}$
$+3p_{2,2}-4p_{5,2}\leqslant 6$ 8.2462 0.272 8.2462 0.296 8.2462 0.310 8.3631 8.2462
288 $-3p_{0,0}-3p_{3,0}+4p_{4,0}+4p_{5,0}-3p_{1,1}-3p_{2,1}$
$+4p_{5,1}-3p_{1,2}-3p_{2,2}+4p_{4,2}\leqslant 8$ 9.8167 0.167 11 0.272 11 0.272 11.7279 9.8167
192 $3p_{1,0}-4p_{5,0}-3p_{0,1}-3p_{2,1}-3p_{3,1}+4p_{4,1}$
$+4p_{5,1}-3p_{1,2}+4p_{4,2}\leqslant 7$ 8.6241 0.178 8.6241 0.188 9.1366 0.221 10 8.6241
192 $3p_{1,0}+3p_{2,0}+3p_{3,0}-4p_{4,0}-3p_{1,1}-3p_{2,1}$
$-3p_{3,1}+4p_{5,1}-3p_{0,2}+4p_{4,2}+4p_{5,2}\leqslant 13$ 14.2154 0.103 15.658 0.249 15.658 0.249 15.8923 14.2154
1728 $\mathcal{I}_{6}^{4}=-3p_{1,0}-6p_{2,0}-6p_{3,0}+4p_{4,0}+8p_{5,0}$
$-6p_{0,1}-3p_{1,1}+4p_{4,1}+8p_{5,1}-3p_{0,2}$ 12.3578 0.024 15.1551 0.199 15.5037 0.218 17.2706 12.3578
$-3p_{2,2}-3p_{3,2}+4p_{4,2}\leqslant 12$
Table 6: In this case, we obtained $5538$ inequalities, among which $38$ are trivial. The rest of the $5500$ inequalities reduce to $10$ inequivalent class as mentioned above. To obtain the inequivalent NCI, we applied the following indistinguishability conditions and symmetry transformations.
$\ p(z|x=6,y)=\frac{3}{4}\left(p(z|x=0,y)+p(z|x=1,y)+p(z|x=2,y)+p(z|x=3,y)\right)-p(z|x=4,y)-p(z|x=5,y)\ \forall y,z$ (66)
$P_{0}\longleftrightarrow P_{1};P_{0}\longleftrightarrow P_{2};P_{0}\longleftrightarrow P_{3};P_{1}\longleftrightarrow P_{2};P_{1}\longleftrightarrow P_{3};P_{2}\longleftrightarrow P_{3};P_{4}\longleftrightarrow P_{5};P_{4}\longleftrightarrow P_{6};P_{5}\longleftrightarrow P_{6}$ (67)
$\ M_{0|y}\longleftrightarrow M_{1|y}\ \forall y;\ \ \ M_{z|0}\longleftrightarrow M_{z|1};\ M_{z|1}\longleftrightarrow M_{z|2};\ M_{z|0}\longleftrightarrow M_{z|2}\ \forall z.$ (68)
orbit size Inequalities $Q_{s}^{2}$ $\omega^{2}_{c}$ $Q_{1}$
48 $-2p_{0,0}+p_{1,0}+p_{2,0}+p_{4,0}\leqslant 1$ 1 0 1
144 $-p_{0,0}+p_{2,0}-p_{0,2}+p_{1,2}\leqslant 1$ 1.4142 0.293 1.4142
144 $2p_{0,0}-p_{2,0}-2p_{4,0}-p_{0,1}+p_{1,1}\leqslant 1$ 1.3371 0.183 1.3638
48 $\mathcal{I}_{7}=p_{0,0}-p_{1,0}+p_{0,1}-p_{4,1}+p_{0,2}-p_{2,2}\leqslant 1$ 1.7321 0.423 1.7321
Table 7: We obtained $384$ nontrivial inequalities that are grouped into $4$ inequivalent classes. It turns out that by rearranging the indistinguishability conditions, we can express the probabilities for four input variables using the other four input variables in the following manner.
$p(z|x=3,y)=p(z|x=1,y)+p(z|x=2,y)-p(z|x=0,y)\ \forall y,z$ (69)
$p(z|x=5,y)=p(z|x=1,y)+p(z|x=4,y)-p(z|x=0,y)\ \forall y,z$ (70)
$p(z|x=6,y)=p(z|x=2,y)+p(z|x=4,y)-p(z|x=0,y)\ \forall y,z$ (71)
$p(z|x=7,y)=p(z|x=1,y)+p(z|x=2,y)+p(z|x=4,y)-2p(z|x=0,y)\ \forall y,z.$ (72)
Interestingly, the symmetries possessed by the preparations $\{P_{x}\}$ exhibit a fascinating connection to the symmetries of a cube. A cube contains $23$ symmetry elements that include centre of symmetry, plane of symmetry and axis of symmetry. The symmetry operations on the preparations along with the corresponding cube symmetries that we apply to find the equivalent NCI are described below.
[…]
$P_{0}\longleftrightarrow P_{7},P_{3}\longleftrightarrow P_{4},P_{2}\longleftrightarrow P_{6},P_{1}\longleftrightarrow P_{5};\text{($\pi$ rotation about an axis dissecting the edges containing $x_{2}x_{6}$ and $x_{1}x_{5}$)}$ (75)
$P_{0}\longleftrightarrow P_{4},P_{3}\longleftrightarrow P_{7},P_{2}\longleftrightarrow P_{5},P_{1}\longleftrightarrow P_{6};\text{($\pi$ rotation about an axis dissecting the edges containing $x_{3}x_{7}$ and $x_{0}x_{4}$)}$ (76)
[…]
$P_{0}\longleftrightarrow P_{6},P_{1}\longleftrightarrow P_{7};\text{(reflection with respect to the diagonal plane containing $x_{2},x_{3},x_{5},x_{4}$)}$ (78)
Note that there are more symmetries, but the ones listed above are enough to group all the NCI into the inequivalent classes. The symmetry transformations implemented on the measurements are given by (62).
Scenario 7
This contextuality scenario involves eight preparations and three measurements, that is, $x\in\{0,1,2,3,4,5,6,7\},y\in\{0,1,2\},z\in\{0,1\}$, satisfying the following indistinguishability conditions:
$\frac{1}{4}\left(P_{0}+P_{1}+P_{6}+P_{7}\right)\sim\frac{1}{4}\left(P_{2}+P_{3}+P_{4}+P_{5}\right)\sim\frac{1}{4}\left(P_{0}+P_{2}+P_{5}+P_{7}\right)\sim\frac{1}{4}\left(P_{1}+P_{3}+P_{4}+P_{6}\right)$ (79)
$\sim\frac{1}{4}\left(P_{0}+P_{3}+P_{4}+P_{7}\right)\sim\frac{1}{4}\left(P_{1}+P_{2}+P_{5}+P_{6}\right)\sim\frac{1}{4}\left(P_{0}+P_{3}+P_{5}+P_{6}\right)\sim\frac{1}{4}\left(P_{1}+P_{2}+P_{4}+P_{7}\right).$
It is worth noting that the indistinguishability conditions in this scenario are not entirely independent. Interestingly, this scenario resembles the $3$-bit parity oblivious multiplexing task, which has been previously explored in Spekkens et al. (2009). In this communication task, the sender possesses a $3$-bit string $x=x_{0}x_{1}x_{2}$ selected uniformly, and the receiver’s goal is to guess the $y$th bit of ${x}$ while adhering to the constraint that all potential parities of the input bits must remain oblivious in the communication. We have found four inequivalent NCI in this scenario, as detailed in Tab. 7. To achieve quantum violations of $\mathcal{I}_{7}$, a specific strategy can be employed wherein $\rho_{x}$ corresponds to qubit states representing the vertices of the cube on the Bloch sphere shown in Tab. 7 within the Bloch sphere, and the reciever’s measurements are $\sigma_{y},\sigma_{x},\sigma_{z}$ for $y=0,1,2$, respectively.
Scenario 8
Until this point, we have considered contextuality scenarios involving indistinguishability conditions on preparations. However, in this particular scenario, we examine eight preparations, with $x\in\{0,1,2,3,4,5,6,7\}$ adhering to the same conditions as specified in (79), and three binary outcome measurements, $y\in\{0,1,2\},z\in\{0,1\}$, which satisfy the indistinguishability condition,
$\frac{1}{3}\{M_{0|0}+M_{0|1}+M_{0|2}\}\sim\frac{1}{3}\{M_{1|0}+M_{1|1}+M_{1|2}\}.$ (80)
The resulting inequivalent NCI can be found in Tab. 8. An important observation here is that, for all the NCI, $Q^{2}_{s}=Q_{1}$ up to machine precision, indicating we have found the exact maximum quantum violations. Moreover, $Q_{1}^{\Pi}=Q_{1}$, implying that projective measurements are sufficient to obtain the maximum quantum violations of all the NCI. A set of qubit states and measurements yielding the maximal violation of $\mathcal{I}_{8}$ is provided in Fig. 2.
orbit size Inequalities $Q_{s}^{2}$ $\omega^{2}_{c}$ $Q_{1}$
48 $-4p_{0,0}+2p_{1,0}+2p_{2,0}+2p_{4,0}-4p_{0,1}+2p_{1,1}+2p_{2,1}+2p_{4,1}\leqslant 3$ 3 0 3
288 $-2p_{0,0}+2p_{1,0}+2p_{2,0}-p_{0,1}+2p_{1,1}\leqslant 3$ 3.3660 0.196 3.3660
288 $5p_{0,0}-3p_{1,0}-2p_{2,0}-2p_{4,0}+2p_{0,1}-2p_{4,1}\leqslant 1$ 1.6458 0.244 1.6458
288 $p_{1,0}-p_{2,0}+p_{4,0}+2p_{0,1}\leqslant 3$ 3.3660 0.196 3.3660
144 $-p_{0,0}+p_{2,0}-p_{0,1}+p_{4,1}\leqslant 1$ 1.3660 0.268 1.3660
144 $-4p_{0,0}+p_{1,0}+p_{4,0}-2p_{0,1}+2p_{2,1}\leqslant 1$ 1.6458 0.244 1.6458
144 $p_{0,0}-3p_{2,0}-4p_{0,1}+2p_{1,1}+2p_{4,1}\leqslant 1$ 1.6458 0.244 1.6458
144 $-3p_{0,0}+p_{1,0}+p_{2,0}+p_{4,0}-2p_{0,1}+p_{1,1}+p_{4,1}\leqslant 1$ 1.3660 0.268 1.3660
48 $\mathcal{I}_{8}=-p_{1,0}+p_{4,0}-2p_{0,1}+p_{2,1}+p_{4,1}\leqslant 1$ 1.5 0.333 1.5
576 $p_{0,0}-2p_{1,0}+p_{2,0}-4p_{0,1}+3p_{2,1}+3p_{4,1}\leqslant 3$ 3.6889 0.256 3.6889
576 $4p_{0,0}-4p_{1,0}-2p_{4,0}-5p_{0,1}+4p_{2,1}+3p_{4,1}\leqslant 3$ 3.9210 0.235 3.9210
Table 8: We obtained $2688$ nontrivial NCI that reduced to $11$ inequivalent classes. For the preparations, we apply the same indistinguishability relations given by (69)-(72) as in Scenario 7. For the measurements, the following relation is imposed.
$p(z=0|x,y=2)=3/2-p(z=0|x,y=1)-p(z=0|x,y=0)\ \forall x.$ (81)
Moreover, we apply the same set of symmetry transformations on both preparations and measurements, as in scenario 7.
Scenario 9
Finally, we consider a scenario where the measurements have three possible outcomes, $z\in\{0,1,2\}$. The indistinguishability conditions in this scenario, consisting of six preparations and two measurements, are given by,
$\frac{1}{2}\left(P_{0}+P_{1}\right)\sim\frac{1}{2}\left(P_{2}+P_{3}\right)\sim\frac{1}{2}\left(P_{4}+P_{5}\right);\quad\frac{1}{2}\{M_{0|0}+M_{0|1}\}\sim\frac{1}{2}\{M_{1|0}+M_{1|1}\},$ (82)
where $x\in\{0,1,2,3,4,5\},y\in\{0,1\}$. The set of inequivalent NCI is given in Tab. 9, and the quantum violation of $\mathcal{I}_{3}$ for qubit systems is illustrated in Fig. 2. For most of the NCI, where $\mathcal{Q}^{\Pi}_{1}<\mathcal{Q}^{2}_{s}$ and the noncontextual bounds are the same as $\mathcal{Q}^{\Pi}_{1}$, any quantum violation certifies nonprojective measurements.
orbit size Inequalities $\mathcal{Q}_{s}^{2}$ $\omega^{2}_{c}$ $\mathcal{Q}_{1}$ $\mathcal{Q}^{\Pi}_{1}$
6 $p^{0}_{4,0}+2p^{0}_{4,1}-p^{1}_{4,0}\leqslant 1$ 1 0 2 0
6 $-2p^{0}_{1,0}-2p^{0}_{1,1}+2p^{1}_{1,0}\leqslant 0$ 0 0 0 0
6 $-2p^{1}_{0,0}-2p^{1}_{1,0}+2p^{1}_{4,0}\leqslant 0$ 0 0 0 0
8 $-2p^{0}_{0,0}-2p^{0}_{1,0}+3p^{0}_{2,0}-2p^{0}_{0,1}-2p^{0}_{1,1}+2p^{0}_{2,1}+$
$2p^{0}_{4,1}+2p^{1}_{0,0}+2p^{1}_{1,0}-p^{1}_{2,0}-2p^{1}_{4,0}\leqslant 1$ 1.4536 0.312 2.8284 1
8 $-2p^{0}_{0,0}-2p^{0}_{1,0}+3p^{0}_{4,0}-2p^{0}_{2,1}+2p^{0}_{4,1}+2p^{1}_{2,0}-p^{1}_{4,0}\leqslant 1$ 1.4536 0.312 2.8284 1
16 $\mathcal{I}_{9}=p^{0}_{1,0}+2p^{0}_{1,1}-2p^{0}_{2,1}-2p^{1}_{0,0}-p^{1}_{1,0}+2p^{1}_{2,0}\leqslant 1$ 1.4536 0.312 2.8284 1
8 $-2p^{0}_{0,0}-2p^{0}_{0,1}-2p^{0}_{1,1}+2p^{0}_{4,1}+2p^{1}_{0,0}-2p^{1}_{4,0}\leqslant 0$ 0.5 0.2 0.8284 0
16 $-2p^{0}_{0,0}-2p^{0}_{1,0}+3p^{0}_{4,0}-2p^{0}_{1,1}+2p^{0}_{4,1}+2p^{1}_{1,0}-p^{1}_{4,0}\leqslant 1$ 1.4536 0.312 2.8284 1
8 $-2p^{0}_{0,0}-2p^{0}_{1,0}+2p^{0}_{4,0}-2p^{0}_{1,1}+2p^{1}_{1,0}-2p^{1}_{4,0}\leqslant 0$ 0.5 0.2 0.8284 0
4 $-2p^{0}_{2,0}-2p^{0}_{0,1}-2p^{0}_{1,1}+2p^{0}_{4,1}+2p^{1}_{2,0}-2p^{1}_{4,0}\leqslant 0$ 0.5 0.2 0.8284 0
2 $-p^{0}_{0,0}-p^{0}_{1,0}+p^{0}_{4,0}-p^{0}_{0,1}-p^{0}_{1,1}+p^{0}_{2,1}+p^{1}_{0,0}+p^{1}_{1,0}-p^{1}_{2,0}-p^{1}_{4,0}\leqslant 0$ 0.25 0.2 0.4142 0
2 $-2p^{0}_{4,0}-2p^{0}_{2,1}-2p^{1}_{0,0}-2p^{1}_{1,0}+2p^{1}_{2,0}+2p^{1}_{4,0}\leqslant 0$ 0.5 0.2 0.8284 0
Table 9: Here, we opt the notation $p_{x,y}^{z}=p(z|x,y)$. We obtained $107$ inequalities, out of which $17$ are trivial. The remaining inequalities are reduced to $12$ inequivalent classes that are listed above. In order to identify the equivalent NCI, the following indistinguishability relations and symmetry transformations are implemented.
$\ p(z|x=3,y)=p(z|x=0,y)+p(z|x=1,y)-p(z|x=2,y)\ \forall y,z$ (83)
$\ p(z|x=5,y)=p(z|x=0,y)+p(z|x=1,y)-p(z|x=4,y)\ \forall y,z$ (84)
$\ p(2|x,y)=1-p(0|x,y)-p(1|x,y)\ \forall x,y$ (85)
$P_{0}\longleftrightarrow P_{1};P_{2}\longleftrightarrow P_{3};P_{4}\longleftrightarrow P_{5}$ (86)
$P_{0}\longleftrightarrow P_{2},P_{1}\longleftrightarrow P_{3};\ P_{0}\longleftrightarrow P_{4},P_{1}\longleftrightarrow P_{5};\ P_{2}\longleftrightarrow P_{4},P_{3}\longleftrightarrow P_{5}$ (87)
$P_{0}\longleftrightarrow P_{3},P_{1}\longleftrightarrow P_{2};\ P_{0}\longleftrightarrow P_{5},P_{1}\longleftrightarrow P_{4};\ P_{2}\longleftrightarrow P_{5},P_{3}\longleftrightarrow P_{4}$ (88)
$\ M_{0|y}\longleftrightarrow M_{1|y}\ \forall y$ (89)
$\ M_{z|0}\longleftrightarrow M_{z|1}\ \forall z.$ (90)
((a)) $\mathcal{I}_{2}=2.6458$ (from Tab. 46)
((b)) $\mathcal{I}_{3}=3.1231$ (from Tab. 3)
((c)) $\mathcal{I}_{5}=2.5$ (from Tab. 62)
((d)) $\mathcal{I}_{6}^{3}=8.2462$ (from Tab. 68)
((e)) $\mathcal{I}_{8}=1.5$ (from Tab. 8)
((f)) $\mathcal{I}_{9}=1.4536$ (from Tab. 9)
Figure 2: The x-z plane of the Bloch sphere is considered to pinpoint the quantum states and measurements that yield the maximum violations of some of the NCI, as determined by the see-saw optimization technique for two-dimensional quantum systems. The symbols $\diamond$ and $\star$ represent the indistinguishable mixed state and the indistinguishable measurement effects in the respective scenario. In figure (2(f)), the length of the Bloch vectors representing $M_{0|0}$ and $M_{1|1}$ is approximately $0.3689$ and the length of the Bloch vectors representing $M_{1|0},M_{0|1}$ is approximately $0.674$.
Figure 2 illustrates the qubit strategies attaining the maximum quantum violations. In many instances, up to machine precision, the maximal quantum violations have been found whenever $\mathcal{Q}_{s}^{d}$ matches with $\mathcal{Q}_{1}$. Of all the NCI, the most resilient one emerges from Scenario 7, with the critical robustness parameter equal to $0.423$. Intriguingly, the best obtained quantum violations do not always stem from cases where the indistinguishable state is the maximally mixed state. Scenario $6$, which features seven preparations and binary outcomes, represents the simplest scenario showcasing variations in quantum violations for different dimensional quantum systems.
Based on the see-saw optimization, a range of NCI exist, violations of which serve as witnesses of the dimension of the quantum systems. For instance, if a quantum violation of $\mathcal{I}_{6}^{1}$ surpasses $7.6833$, then the dimension of the quantum systems must be at least three. Moreover, Scenario $9$ provides several NCI whose violations using qubits certify three outcome non-projective measurements. In fact, non-projective measurements are necessary for achieving quantum violations of all the NCI in this contextuality scenario.
V Applications of newly found NCI
In this section, we delve into some of the interesting applications of newly found NCI.
Certification of non-projective measurements
The study by Chaturvedi et al. Chaturvedi et al. (2021b) points out that the maximum value of an NCI can always be achieved using projective measurement, where no indistinguishability conditions on measurements are imposed. Consequently, certification of non-projective measurements through the violation of NCIs can only occur with nontrivial indistinguishable conditions on measurements. To accomplish this, it is necessary to establish an upper bound on the expression of NCIs when measurements are restricted to be projective for arbitrary dimensional quantum states and measurements. As outlined in Chaturvedi et al. (2021a), the semi-definite hierarchy can be modified to obtain upper bounds when the measurements are projective. Let us denote these upper bounds by $\mathcal{Q}_{1}^{\Pi}$.
We have implemented this optimization to obtain $\mathcal{Q}_{1}^{\Pi}$ in the last two scenarios under consideration, both featuring nontrivial indistinguishability conditions for measurements. Among these scenarios, it was found that in scenario VIII, $\mathcal{Q}_{1}$ is the same $\mathcal{Q}_{1}^{\Pi}$ for all NCIs, suggesting this scenario is unable to certify non-projective measurements. However, in scenario IX, the $\mathcal{Q}_{1}^{\Pi}$ values are lower than $\mathcal{Q}_{s}^{2}$ for all NCIs whenever a quantum violation occurs. These precise values are documented in the final column of Table 9. Notably, the upper bounds for projective measurement coincide with the noncontextual values for all NCIs, indicating that any violation of these NCIs implies unequivocally that the measurements are non-projective.
Noncontextuality inequalities as dimension witnesses
As a consequence of our extensive investigations, we reveal a noteworthy aspect of the interplay between quantum preparation contextuality and quantum Hilbert space dimension. In particular, we find that seven of ten noncontextuality inequalities in Scenario 6, Table 68, double as dimension witnesses for Hilbert space dimension $d>2$. Specifically, we implemented a novel hierarchy of SDP relaxations. In particular, the novel scheme employs the scheme described in Chaturvedi et al. (2021a) over the basis generated from the Navascués-Vértesi method for bounding finite-dimensional quantum correlations and retrieve tight upper bounds $Q^{2}_{UB}=Q^{2}_{s}$ with level $3$ for inequalities in Table 68. Notably, the lower bounds obtained from the see-saw for dimension $d>2$ violate the upper bounds for seven noncontextuality inequalities, thereby forming here-to-unknown non-trivial dimension witnesses. It is worth mentioning that $\mathcal{I}_{6}^{2}$ in Tab.68 is violated by qutrit systems, while qubit states fail to produce any violation. Moreover, without the operational equivalences, the dimension restriction fails to yield non-trivial bounds on the inequalities; the operational equivalences are necessary to witness the Hilbert dimension $d>2$ with these noncontextuality inequalities.
Randomness certification
In this section, we demonstrate that novel instances of quantum contextuality found in this article can be used for semi-device-independent quantum randomness certification with operational equivalences. To exemplify this observation, we consider the inequality $\mathcal{I}_{7}$ in Table 7 found in Scenario 7. In particular, We evaluate the extractable randomness in Bob’s output associated with the violation of the noncontextuality inequality $\mathcal{I}_{7}\in(1,1.7321]$. Let us suppose that the parties decide to extract randomness from Bob’s first measurement $y=0$ performed on Alice’s first input $x=0$. The certified randomness in this case is gauged min-entropy $H_{min}=-\log_{2}{p^{*}}$ where $p^{*}=\max{p_{0,0},1-p_{0,0}}$ given the value of $\mathcal{I}_{7}$. We use the hierarchy of SDP programs Chaturvedi et al. (2021a) to retrieve upper bounds on $p^{*}$, which translate to lower bounds on the $H_{min}$. In FIG. 3 we plot the lower bounds on certified randomness $H_{min}$ obtained from level 3 of the SDP hierarchy against the violation of $\mathcal{I}_{7}$. We find that non-zero randomness can be certified in the range $\mathcal{I}_{7}\in[1.52,1.7321]$ with maximum $0.34147$ coinciding with the maximum attainable violation $\mathcal{I}_{7}=Q^{2}_{s}=Q_{1}$ (up to machine precision). We note here that the choice of $x=0,y=0$ turns out be optimal and equivalent to $x=0,y=1$ and $x=0,y=2$.
Figure 3: Randomness ($H_{min}$) as a function of $\mathcal{I}_{7}\in[1.5,1.7321]$ from Table 7.
Quantum advantage in oblivious communication
Oblivious transfer is a crucial task in information theory with myriad applications in cryptography. In Saha et al. (2019), the oblivious transfer task has been generalized, and it is shown that any quantum advantage in such tasks, namely, oblivious communication task, implies preparation contextuality. Moreover, quantum violations of preparation NCIs in prepare-and-measure scenarios directly translate into quantum advantages in oblivious communication tasks. Here, the oblivious condition is associated with the indistinguishable condition on the preparations, and the expression of the respective NCI corresponds to the figure of merit of that task. Consequently, any quantum violation of newly discovered preparation NCIs can be interpreted as a quantum advantage in an oblivious communication task.
In this regard, an interesting fact emerges from the discussion after Table 46. The optimal classical encoding strategy for the sender in the oblivious communication task with respect to $\mathcal{I}_{2}$ must be probabilistic for saturating the noncontextual bound $2$, and the maximum value of $\mathcal{I}_{2}$ for deterministic encoding strategies is $1$.
VI Conclusions
Deriving a set of empirical criteria applicable to any operational theory that satisfies the generalized notion of contextuality is an arduous task of both foundational and operational significance. The conventional method of extracting facet inequalities from the pertinent noncontextual polytope is computationally demanding due to the polynomial growth in the dimension of the polytope describing the preparations with the number of measurements. This study introduces an innovative approach for constructing a polytope that encompasses the actual noncontextual polytope while ensuring that the complexity of the method remains minimal. The facet inequalities resulting from the intersection of this extended polytope with the normalization polytope constitute necessary conditions for noncontextuality (see Fig. 1). To validate this methodology, it is applied to nine simple scenarios comprising four to nine preparations and two to three measurements, leading to the respective sets of inequalities. It is worth noting that there could potentially be further symmetries present in certain contextuality scenarios that haven’t been accounted for during the process of deriving the inequivalent classes of NCI. These unexplored symmetries might lead to a reduction in the number of distinct sets of NCI.
Of all the NCI, the most resilient one emerges from Scenario 7, with the critical robustness parameter equal to $0.423$. In many instances, up to machine precision, the maximal quantum violations have been found whenever $\mathcal{Q}_{s}^{d}$ matches with $\mathcal{Q}_{1}$. Figure 2 illustrates the qubit strategies attaining the maximum quantum violations. Intriguingly, the best obtained quantum violations do not always stem from cases where the indistinguishable state is the maximally mixed state. Scenario $6$, which features seven preparations and binary outcomes, represents the simplest scenario showcasing variations in quantum violations for different dimensional quantum systems. A set of NCI exist, violations of which serve as witnesses of the dimension of the quantum systems. Moreover, Scenario $9$ provides several NCI whose violations using qubits certify three outcome non-projective measurements. In fact, non-projective measurements are necessary for achieving quantum violations of all the NCI in this contextuality scenario.
This study has focused on a set of indistinguishability conditions regarding preparations (or measurements) to render them indistinguishable from each other. Nonetheless, it’s possible to consider scenarios with more than one set of indistinguishability conditions, each corresponding to convex decompositions of mixed preparations (or measurements) Chaturvedi et al. (2021a). Extending our method to cover such scenarios is a straightforward task that could be explored more thoroughly in future research.